OpenStudy (anonymous):

3^x2+4x=1/27 (the x2+4x is all raised to the 3)

OpenStudy (anonymous):

3^( x²+4x) = 3^ (-3) can you continue?

OpenStudy (anonymous):

No i'm lost... where did the 3^(-3) come from on the right side?

OpenStudy (anonymous):

1/ 27 = 1/ 3³

OpenStudy (anonymous):

okay so do I just solve (x2+4x=-3) now?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

Thank you